A force of 230N exist between two identical point charges of 56uC. Calculate the distance of the two point charge.


1
Expert's answer
2022-03-30T14:09:44-0400
"F=k\frac{q_1q_2}{r^2 }"

"r=\sqrt{kq^2/F}\\
=\sqrt{9*10^9*(56*10^{-6})^2/230}=0.35\:\rm m"